# Large Language Models, scientific knowledge and factuality: A systematic analysis in antibiotic discovery

### v2

@@ Abstract
Inferring over and extracting information from Large Language Models (LLMs) trained on a large corpus of scientific literature can potentially drive a new era in biomedical research, reducing the barriers for accessing existing medical evidence. This work examines the potential of LLMs for dialoguing with biomedical background knowledge, using the context of antibiotic discovery. The systematic analysis is applied to ten state-of-the-art models, from models specialised on biomedical scientific corpora to general models such as ChatGPT, GPT-4 and Llama 2 in two prompting-based tasks: chemical compound definition generation and chemical compound-fungus relation determination. The work provides a systematic assessment on the ability of LLMs to encode and express these relations, verifying for fluency, prompt-alignment, semantic coherence, factual knowledge and specificity of generated responses. Results show that while recent models have improved in fluency, factual accuracy is still low and models are biased towards over-represented entities. The ability of LLMs to serve as biomedical knowledge bases is questioned, and the need for additional systematic evaluation frameworks is highlighted. The best performing GPT-4 produced a factual definition for 70% of chemical compounds and 43.6% factual relations to fungi, whereas the best open source model BioGPT-large 30% of the compounds and 30% of the relations for the best-performing prompt. The results show that while LLMs are currently not fit for purpose to be used as biomedical factual knowledge bases, there is a promising emerging property in the direction of factuality as the models become domain specialised, scale-up in size and level of human feedback.
@@ 1 Introduction
Recently, artificial intelligence (AI) algorithms have significantly facilitated the acceleration of antibiotic discovery through high-throughput drug screening, with a particular emphasis on the identification of novel antibiotic compounds [1]. Deep learning (DL) methods have proven effective in isolating small molecule antibiotics exhibiting potent activity against a diverse array of bacterial pathogens [2]. This advancement in machine learning (ML) for drug discovery has been bolstered by the widespread accessibility of datasets emerging from empirical analyses. Notably, the discovery of novel compounds primarily relies on protein structure prediction and modeling of antibiotic targets. Moreover, since the knowledge related to antibiotic discovery is communicated as text in the scientific literature, patents and clinical trials, this creates an opportunity for Large Language Models (LLMs), especially Generative Pre-trained Transformers (GPTs), to serve as proxy for background scientific knowledge, reducing the barriers for integrating and mapping the landscape of available evidence. Eliciting information from LLMs trained on large corpora of scientific, biomedical literature could potentially support the antibiotic discovery process.
Large Language Models have evolved in the past two years significantly influencing natural language processing (NLP) with a variable level of adoption in different specific domains [3, 4]. GPT-1 (Generative Pre-trained Transformer) [5], the first LLM trained using self-supervised pre-training, predicts the next word in a sequence given the previous words, performing this task unidirectionally from left to right. This enables GPT-1 to learn the dependency relations between words and generate relevant and contextual content. GPT-2 [6] was trained with a larger architecture on a larger and more diverse corpus than GPT-1, with 1.5 billion parameters, thus demonstrating better generalisation properties.
GPT-3 [7] with 175 billion parameters represents a phase transition in the evolutionary arch of language models, where the in-context learning properties were noticeably present. Subsequently, InstructGPT [8] extended this foundation, using reinforcement learning from human feedback (RLHF) as a training strategy, leveraging on the ability of the pre-training model to re-align at instruction/intensional level. ChatGPT [9] follows the same architectural principles of InstructGPT, using GPT-3.5 (both larger corpora, and human-feedback prompting datasets). Despite having significantly less domain-specific biomedical text, it can still deliver the ability to encode and articulate more specialised discourse, for example, achieving a score equivalent to a third-year medical student [10].
@@
Despite their comprehensive linguistic interpretation properties, LLMs have inherent operational properties, most notably hallucinations [11, 12], i.e. divergence from factual knowledge and spurious inference, despite its fluency and syntactic correctness [13, 14]. The recently released GPT-4 is not free from this limitation [15]. Hallucinations tend to surface prominently in domain-specific discourse, particularly in biomedical sciences. This elicits the need to establish the ability of these models to encode and preserve factual knowledge as well as establishing methodologies for the critical evaluation of its representation properties [16, 17, 18].
In a range of benchmarks assessing helpfulness and safety, the one of the latest open-source LLMs, Llama 2, typically outperforms current open-source alternatives. Llama 2 was pretrained on publicly available online data sources (on 2 trillion tokens) and its fine-tuned models have been trained on over 1 million human annotations. As the Llama2’s authors [19] emphasized, testing conducted to date has been in English and has not — and could not — cover all scenarios. None of the scenarios tested so far included the use of the model for the purpose discussed in this study. We evaluated variants of this model with 7B, 13B, and 70B parameters.
Figure 1: The overview of the study. Investigation of the ability of LLMs to encode and access biological factual/relational knowledge. More specifically, a highly relevant domain of application was selected, which is motivated by antibiotic discovery.
@@
In this work we systematically assess the biological relational knowledge encoded in state-of-the-art LLMs, namely: GPT-2, GPT-3, GPT-neo, Bloom, BioGPT, BioGPT-Large, Galactica, ChatGPT, GPT-4 and variants of Llama 2 with 7B, 13B, and 70B parameters. We focus on systematically determining the ability of LLMs to capture fundamental entities (fungi, chemicals and antibiotic properties), their relations and supporting facts, which are fundamental for delivering inference in the context of downstream biomedical inference (Figure 1). For instance, we aim to answer the question on whether LLMs can faithfully capture biological domain-specific factual knowledge such as:
  • “Aspergillus fumigatus produces festuclavine"
  • “Cycloclavine is biosynthesised by Aspergillus japonicus"
  • “Enfumafungin demonstrates antibiotic activity against several bacterial, yeast and fungal species"
  • “Enfumafungin and ergokonin A are isolated from Hormonema and Trichoderma species and demonstrate antibiotic activity against several bacterial, yeast and fungal species"
@@ 1. RQ1
    Do large language models encode biomedical domain knowledge at entity level (e.g. fungi, chemical, antibiotic properties) and at a relational level?
  2. RQ2
@@
    Are there significant differences between different models in their ability of encoding domain knowledge?
  3. RQ3
@@
    Which model performs better in faithfully encoding biological relations?
  4. RQ4
    How do we evaluate the limitations of LLMs in factual knowledge extraction?
This work targets the following contributions:
  • Introduction of a comprehensive framework tailored to systematically assess LLMs capacity to encapsulate domain-specific factual knowledge, focusing on target entities and their interrelations within the biomedical sphere. This framework is a direct response to the labor-intensive nature of text generation evaluation and aims at reducing this burden
  • We provide a qualitative evaluation and comparison of the state-of-the-art LLMs in the context of biomedical inference and factual knowledge extraction. We broaden the scope of existing research from other fields, catering to the often-expressed demand for more comprehensive evaluations of LLMs across diverse disciplines.
  • We explore the extent to which these models rely on memorization versus genuine comprehension of the intricate symbolic relationships between biomedical terms
  • Finally, we present a compelling case and substantial proof supporting the necessity for evaluation frameworks. These frameworks are crucial for identifying the limitations of LLMs and ensuring the effective evaluation of the text they generate.
@@ 4. RQ4
This study is organized as follows: first, we define our contribution and link it with the existing frameworks and other investigations within the biomedical domain. Then, in the Methodology section, we present a framework for the qualitative evaluation of the biomedical knowledge embedded in LLMs, followed by a detailed description of the methods. In the Results section we apply the proposed framework to ten LLMs and evaluate them for ten chemical compounds with a detailed description of the generated responses. We then discuss our findings in the broader context of reported LLMs’ limitations. We review related work in the field of LLMs and their knowledge representation properties. Finally, we conclude the paper with a summary of the contributions and with future perspectives on the role of LLMs in biomedical inference.
The workflow of the analysis is depicted in Figure 2, and the proposed evaluation framework in Figure 3.
2 Motivational scenario
Antimicrobial resistance in clinically relevant bacteria compromises the effectiveness of existing antibiotics and call for the discovery of new antibiotic substances. Antibiotic development is a slow, costly and failure-prone process that can take over a decade and costs can reach the scale of hundreds of millions of dollars [20]. In a study of nearly 186,000 clinical trials spanning over 21,000 compounds, the probability of success of new drugs treating infectious diseases was found to be 25.2% [21]. However since 2017, only eight new antibiotics have been approved, including one for the treatment of multidrug-resistant tuberculosis (2017-2021) [22]. Furthermore, antibiotics in clinical trials are generally analogues of existing drugs for which antimicrobial resistance mechanisms have already emerged [23], further highlighting the need for new approaches to antibiotic discovery.
Some organisms, such as fungi or bacteria, naturally produce a variety of secondary metabolites (also called natural products) associated with many biological activities. Penicillin, a beta-lactam antibiotic, is a historical example of fungal metabolites with therapeutic properties. In order to effectively exploit the pool of potential sources of antibiotics, avoiding dead ends and rediscovery, it is essential to identify the secondary metabolites known to be produced by these organisms and which substances have already shown antibiotic activity. This knowledge is mainly communicated in the scientific literature, the size and complexity of which makes it difficult to explore it systematically and at scale.
3 Methodology
In this section we describe the workflow of our analysis (depicted in Figure 2), which consists of the following steps:
  • extraction and curation of a dataset with biological relations
  • definition of text generation tasks for LLMs
  • selection of the target LLMs to be included in the evaluation
  • development of an evaluation framework for the efficient qualitative analysis of generated output, including prompt definition, prompt engineering and evaluation metrics
@@ 3 Methodology
The above steps allowed us to provide the first comparative analysis of LLMs in the context of tasks related to fungi and antibiotics and is transportable to other biological and biomedical relations. The code and dataset used in the study are available at github [^1] https://github.com/digital-ECMT/LLM_factuality_evaluation_framework.git.
@@
Figure 2: The workflow of the performed analysis. $n_{relations}$ 23 selected chemical compound-fungus pairs; $n_{entity1}$ 10 unique chemical compounds included in those pairs; In prompt engineering: blue - entity 1, orange - added context. Qualitative evaluation performed according the the framework depicted in Fig.3.

### v3

# Large Language Models, scientific knowledge and factuality: A framework to streamline human expert evaluation

@@ Abstract
The paper introduces a framework for the evaluation of the encoding of factual scientific knowledge, designed to streamline the manual evaluation process typically conducted by domain experts. Inferring over and extracting information from Large Language Models (LLMs) trained on a large corpus of scientific literature can potentially define a step change in biomedical discovery, reducing the barriers for accessing and integrating existing medical evidence. This work explores the potential of LLMs for dialoguing with biomedical background knowledge, using the context of antibiotic discovery. The framework involves of three evaluation steps, each assessing different aspects sequentially: fluency, prompt alignment, semantic coherence, factual knowledge, and specificity of the generated responses. By splitting these tasks between non-experts and experts, the framework reduces the effort required from the latter. The work provides a systematic assessment on the ability of eleven state-of-the-art models LLMs, including ChatGPT, GPT-4 and Llama 2, in two prompting-based tasks: chemical compound definition generation and chemical compound-fungus relation determination. Although recent models have improved in fluency, factual accuracy is still low and models are biased towards over-represented entities. The ability of LLMs to serve as biomedical knowledge bases is questioned, and the need for additional systematic evaluation frameworks is highlighted. While LLMs are currently not fit for purpose to be used as biomedical factual knowledge bases in a zero-shot setting, there is a promising emerging property in the direction of factuality as the models become domain specialised, scale-up in size and level of human feedback.
@@ 1 Introduction
Despite comprehensive linguistic interpretation properties, LLMs have inherent operational limitations, most notably hallucinations Bender et al. (2021); Ji et al. (2023a), i.e. divergence from factual knowledge and spurious inference, despite its fluency and syntactic correctness Mahowald et al. (2023); Weidinger et al. (2022). This elicits the need to establish the ability of these models to encode and preserve factual knowledge as well as to define methodologies for the critical evaluation of its representation properties Wysocki et al. (2023a); Jullien et al. (2022); Rozanova et al. (2023).
In this work we introduce an evaluation framework, which is designed to streamline the manual evaluation process typically conducted by domain experts. Despite the advancement and implementation of automated evaluation methods Li et al. (2024), the necessity for manual, human-led verification persists Bavaresco et al. (2024). Users engaged in the biomedical field are likely to continue performing assessments of models within the context of new tasks, using examples they are well-acquainted with to ensure the model’s accuracy and reliability. The framework, therefore, is not meant to replace human evaluation but to guide users on key aspects to consider when conducting their assessments, acting as a roadmap for thorough and effective manual evaluation. Being transferable to other biomedical domains, it enables the recognition of, i.a., biases in LLM outputs towards prevalent topics, challenges in identifying rare entities, deviations from intended contexts, and significant performance variability influenced by prompt design.
@@
Using the framework, we systematically assess the biological relational knowledge encoded in a broad spectrum of state-of-the-art LLMs, namely: GPT-2, GPT-3, GPT-neo, Bloom, BioGPT, BioGPT-Large, ChatGPT, GPT-4 and variants of Llama 2 with 7B, 13B, and 70B parameters. Recently, artificial intelligence (AI) algorithms have significantly facilitated the acceleration of antibiotic discovery through high-throughput drug screening, with a particular emphasis on the identification of novel antibiotic compounds Jablonka et al. (2024); Guo et al. (2023); Torres and de la Fuente-Nunez (2019); Lluka and Stokes (2023); Ruiz Puentes et al. (2022); David et al. (2021); Melo et al. (2021). Motivated by this progress, we focus on systematically determining the ability of LLMs to capture fundamental entities (fungi, chemicals and antibiotic properties), their relations and supporting facts, which are fundamental for delivering inference in the context of downstream biomedical inference. For instance, we aim to answer the question on whether LLMs can faithfully capture biological domain-specific factual knowledge such as: “Aspergillus fumigatus produces festuclavine"; “Cycloclavine is biosynthesised by Aspergillus japonicus".
@@
RQ1: Do large language models encode biomedical domain knowledge at entity level (e.g. fungi, chemical, antibiotic properties) and at a relational level?
@@
RQ2: Are there significant differences between different models in their ability of encoding domain knowledge?
@@
RQ3: Which model performs better in faithfully encoding biological relations?
RQ4: How do we evaluate the limitations of LLMs in factual knowledge extraction?
@@
This study is organised as follows: first, we define our contribution and link it with the existing frameworks and other investigations within the biomedical domain. Then, within the Materials and methods section, we present a framework for the qualitative evaluation of the biomedical knowledge embedded in LLMs. In the Results section, we apply the proposed framework to eleven LLMs and evaluate them for ten chemical compounds with a detailed description of the generated responses. We then discuss our findings in the broader context of reported LLMs’ limitations. Finally, we conclude the paper with a summary of the contributions and with future perspectives on the role of LLMs in biomedical inference.
2 Related Work
LLMs achieve state-of-the-art performance in multiple Natural Language Inference and Understanding tasks, frequently outperforming previous baselines in few-shot settings Brown et al. (2020). The knowledge extracted by prompting these models has been widely discussed in the context of its factuality Ji et al. (2023b); Bavaresco et al. (2024). Since the first demonstration by Petroni et al. Petroni et al. (2019) on Masked Language Models, several studies have examined the mechanisms and biases underlying these outputs, questioning their applicability beyond their potential as knowledge bases.
Howard et al. Howard et al. (2023) examined the implications of generative AI for antimicrobial advisory for addressing infection questions and outlined the limitations of implementing ChatGPT for interventions due to deficits in situational awareness, inference, and consistency. Li et al. Li et al. (2023) reported that ChatGPT performs moderately or poorly in several biomedical tests and is unreliable in actual clinical use. Furthermore, integrating LLMs into clinical practice is associated with numerous obstacles, such as deficits in situational awareness, lack of reasoning control mechanisms, and consistency Wang et al. (2020). While the use of natural language processing in healthcare is not novel, LLMs have elicited intense debate regarding its potential opportunities and challenges in healthcare Nori et al. (2023).
This paper builds upon findings from several key studies to address the limitations and potentials of LLMs in encoding scientific domain knowledge. Zhao et al. Zhao et al. (2021), Kassner, Krojer, and Schutze Kassner et al. (2020) and Wysocki et al. Wysocki et al. (2023a) emphasise the inconsistency of language models and their tendency to replicate prevalent responses, a form of co-occurrence bias also noted in LLMs by Kandpal et al. Kandpal et al. (2023) and Kang and Choi Kang and Choi (2023). This tendency undermines the models’ ability to generalise, a concern supported by Razeghi et al. Razeghi et al. (2022), who report an accuracy disparity tied to the frequency of terms in training data. Biderman et al. Biderman et al. (2023), Power et al. Power et al. (2022) and Tirumala et al. Tirumala et al. (2022) explore delicate balance between memorisation’s beneficial role in preventing the hallucinations, versus outputting entire sequences from their training data verbatim. This is a key safety concern, particularly in preventing the unintended disclosure of sensitive or proprietary information from the training data Carlini et al. (2021). Wang et al. Wang et al. and Delmas, Wysocka, and Freitas Delmas et al. (2023) remark on LLMs’ underperformance when compared to supervised settings, positing retrieval-augmentation as a potential remedy for knowledge-dense tasks. Together, these studies highlight the need for strategies to refine LLMs’ bias and factuality, enhancing their application in biomedical domain.
The essential requirement for the biomedical utility of LLMs is the factual accuracy of the generated text. Standard N-Gram matching metrics, including BLEU Papineni et al. (2002), ROUGE Lin (2004), and token-level F1, have been found to poorly correlate with factual consistency Maynez et al. (2020a); Honovich et al. (2021). Given the cumbersome nature of human evaluation for factuality, there is a rising inclination towards employing LLMs for automating the evaluation of model output. Several studies suggest that LLMs can effectively serve as evaluators, often aligning well with human assessments Liu et al. (2023); Zheng et al. (2023); Chen et al. (2023); Törnberg (2023); Huang et al. (2024); Naismith et al. (2023); Gilardi et al. (2023); Kocmi and Federmann (2023); Verga et al. (2024), though there are some exceptions Wang et al. (2023); Wu and Aji (2023); Hada et al. (2024); Pavlovic and Poesio (2024). In contrast, other research highlights notable deficiencies in the performance of LLMs as evaluators Koo et al. (2023); Zeng et al. (2024); Baris Schlicht et al. (2024), and some studies fail to benchmark LLMs against human judgments Jiang et al. (2023); Landwehr et al. (2023). For summary LLM-based evaluation methodologies, we refer to Li et al. (2024). Bavaresco et al. conclude that LLMs are not yet equipped to systematically replace human judges in NLP, reinforcing our argument for the need for human evaluation Bavaresco et al. (2024).
Existing frameworks for evaluating the factuality of LLMs highlight various approaches and their effectiveness. Luo et al. proposes a systematic framework using diverse, high-coverage questions generated from Knowledge Graphs Luo et al. (2023). Hendrycks et al. measures how well text models learn and apply knowledge encountered during pretraining, assessing language understanding across 57 subjects of varying difficulty Hendrycks et al. (2020). The FRANK survey evaluates the faithfulness metrics for summarisation and compares the correlations of these metrics with human judgments, introducing a typology of errors for factual consistency Pagnoni et al. (2021). Rashkin et al. suggests that binary labeling is more beneficial for practical applications where filtering out unfaithful predictions is necessary, aligning with recommendations for human evaluation of attribution in text generation Rashkin et al. (2021). Sun et al. provides a comprehensive study on the trustworthiness of LLMs, establishing benchmarks and principles for different dimensions of trustworthiness Sun et al. (2024).
In the bioinformatics domain, Yin et al. provides a thorough analysis of the limitations of LLMs in complex bioinformatics tasks Yin et al. (2024). Piccolo et al. evaluates LLMs in bioinformatics programming, focusing on their ability to handle domain-specific coding tasks Piccolo et al. (2023). Park et al. conducts a comparative evaluation of LLMs for automatically extracting knowledge from scientific literature to understand protein interactions and pathway knowledge Park et al. (2023). Our work introduces a unique evaluation framework that not only rigorously tests LLMs’ ability to encode and apply biomedical knowledge in antibiotic discovery but also significantly reduces the effort required from human experts. This focus on minimizing human evaluation sets our approach apart from previous studies, providing new insights and actionable recommendations to improve factual accuracy and domain-specific reliability.
3 Materials and methods
Figure 1: The framework to streamline human expert evaluation of LLMs and the encoding of factual scientific knowledge in the Large Language Models, in the context of extracting biological relations. Entity 1 stands for chemical compound name. The entity 2 stands for fungus name. Fluency, prompt-alignment, and semantic coherence are assessed in STEP 1 by a non-expert (within the target domain). Then the domain expert evaluates the factuality in STEP 2 for Task 1. For Task 2, the factuality of the generated entity 2 (STEP 2A) is verified before the entire description is assessed (STEP 2B). Outputs that do not pass STEP 2 are classified as hallucinations. The specificity is evaluated in STEP 3.
3.1 A framework to streamline human expert evaluation of LLMs on the encoding of factual scientific knowledge
Our evaluation framework, depicted in Fig. 1, is designed to pragmatically assess the biomedical utility of LLMs by closely mirroring the human evaluation process from which it is directly derived. We propose a methodology that substantially reduces the time commitment required from experts, acknowledging the high value and limited availability of their expertise. Our approach formalises a system where non-experts, guided by specific criteria, can filter out outputs that do not require specialised biomedical knowledge, relying instead on their language proficiency. As a result, the volume of outputs that need expert evaluation is significantly reduced.
Existing human evaluation frameworks highlight distinct dimensions of this process: emphasising the impact of cognitive and utility biases on evaluation, particularly with perception-based metrics for truthfulness Elangovan et al. (2024); examining whether LLMs can effectively assist bioinformatics experts Chen and Deng (2023); proposing a framework for model responses along multiple dimensions, including factuality, comprehension, reasoning, potential harm, and bias Singhal et al. (2023).
Our framework complements these by aiming to reduce the effort required from human experts when their involvement is necessary, an area which is still unexplored by existing frameworks. To the best of our knowledge, no existing framework optimises the evaluation process by composing non-expert and expert assessment phases. We argue that as LLMs are increasingly used in biomedical research, manual human evaluation will always be an integral part of model assessment.
The framework consists of three steps: 1) a non-expert evaluates whether the LLM produces text that is fluent, aligned with the prompt, and semantically coherent. If these criteria are met, the output is then passed on to an expert for further evaluation; 2) the expert verifies the factual accuracy of the output; 3) the expert assesses whether the output is generic or specific to the prompt. Each criterion is assessed qualitatively, receiving a binary score of either 0 or 1. A detailed description of the metrics is provided below.
Fluency evaluates the quality of the generated text, considering two sub-criteria: syntactic correctness and style Kann et al. (2018); Mutton et al. (2007). Syntactic correctness assesses whether a sentence conforms to grammar. Fluency is assessed first due to its simplicity, allowing non-experts to perform the evaluation effectively.
Prompt-alignment refers to the relation of relevance and pragmatic coherence between the prompt and the generated text, whether the generated text is consistent with the input prompt and is related to the entities and context of the prompt Webson and Pavlick (2021); Raj et al. (2022). In other words, it checks whether the model accurately answers the question without drifting into irrelevant responses - an aspect that non-experts can effectively identify.
Semantic coherence consists of two levels of analysis: (i) intra-sentence assesses whether the sentence is relating surface terms to expected argument types for a given predicate and (ii) inter-sentence: assessing whether adjacent sentences have a coherent discourse relation Ke et al. (2022).
Factuality is distinct from the meaning of a sentence which is conveyed by its semantic coherence. A sentence generated by a model is true if it aligns with a factual statement in the scientific literature, a database, or any other trusted source Maynez et al. (2020a).
Specificity evaluates the alignment between the prompt and the answer regarding their level of abstraction. While some answers can be factually correct, they may not answer to the level of abstraction required by the prompt Maynez et al. (2020a).
While auto-regressive language models are designed to deliver fluency, prompt-alignment and semantic coherence, text generation during decoding can lead to repetitive, incoherent or meaningless outputs Holtzman et al. (2020). As they are complementary, the fluency, prompt-alignment and semantic coherence were assessed in one step of the evaluation process (Fig. 1, STEP 1). Moreover, the probabilistic nature of the next token prediction task also contrasts with the notion of factuality and large language models are prone to hallucination Maynez et al. (2020b); Ji et al. (2023b). Hallucinations are factually incorrect statements which can be difficult to be localised by non-experts due to their expression in a fluent and semantically coherent text Varshney et al. (2023); Curran et al. (2023). In the context of this study, major hallucinations can be associated with the compound description or relationships with fungi, while minor hallucinations may be associated with the date of first isolation for instance. In our pragmatic approach, any output that reaches the factuality check (fulfills STEP 1) but fails to pass is considered a hallucination.
@@ 3 Materials and methods
This framework encompasses prompt definition, prompt engineering, and evaluation metrics. Then the analysis includes the extraction and curation of a dataset with biological relations, the definition of two text generation tasks for LLMs, and the selection of the target LLMs for evaluation (Fig. 2). These steps allowed us to provide the first comparative analysis of LLMs in the context of tasks related to fungi and antibiotics and is transportable to other relations in bioscience, such as "A is produced by B," "A inhibits B","A activates B," etc. The code and dataset used in the study are available on GitHub [^1] https://github.com/digital-ECMT/LLM_factuality_evaluation_framework.git.
@@
Figure 2: The workflow of the performed analysis. $N$ whole dataset with 246 selected biological relations of chemical compound-fungus pairs; $n_{relations}$ 23 selected chemical compound-fungus pairs; $n_{entity1}$ 10 unique chemical compounds included in those pairs; In prompt engineering: blue - entity 1, orange - added context. Qualitative evaluation performed according the framework depicted in Fig. 1.